Solve the problem.
Find the length of a rectangular lot with a perimeter of 72 m if the length is 8 m more than the width

Found 2 solutions by rfer, checkley77:
Answer by rfer(16322)

2x+2(x+8)=72
2x+2x+16=72
4x=56
x=14 m
x+8=22 m

Answer by checkley77(12844)

L=W+8
2L+2W=72
2(W+8)+2W=72
2W+16+2W=72
4W=72-16
4W=56
W=56/4
W=14 ANS. FOR THE WIDTH.
L=14+8=22 ANS. FOR THE LENGTH.
PROOF:
2*22+2*14=72
44+28=72
72=72